OpenStudy (anonymous):

Can you please show the steps? My Answer was "No Solution" 5x-[7-(2x-1)]=3(x-5)+4(x+3)

OpenStudy (anonymous):

5x-[7-(2x-1)]=3(x-5)+4(x+3) 5x-[7-2x+1]=3x-15+4x+12 5x-7+2x-1=3x-15+4x+12 5x+2x-3x-4x=-15+12+7+1 it's not no solution

OpenStudy (anonymous):

wait i stand corrected it's no solution

OpenStudy (anonymous):

so is it no solution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh ok